if 7.989 g of CuNO3 is dissolved in water to make a 0.0370 M solution, what is the volume of the solution?
Answer #1

Molar mass of CuNO3 = 125.5 gm

Number of moles= mass/ molar mass
    = 7.989 / 125.5
    =0.0637 mol

Molarity = number of moles /volume
0.370 = 0.0637 / volume
volume = 0.172 L or 172 mL
Answer #2

Molar mass of CuNO3 = 125.55 gm

moles of CuNO3 = 7.989/125.55 moles = 0.06363 mole

Molarity = mole/volume

volume = mole/molarity = 0.06363/0.370 = 0.1719 litre ...answer
Answer #3

molarity = \frac{no. of moles}{volume(Litre)}

givenmass molecularmass no.ofmoles

molecular mass of CuNO3 = 126 g

given mass of CuNO3 = 7.989 g

no. of moles = 7.989 / 126 = 0.0634

from the first equation we can write,

Volume = no. of moles / molarity

           = 0.0634 / 0.0370 = 1.71 L